### v1

Abstract
We develop and test high-order methods for integration on surface point clouds. The task of integrating a function on a surface arises in a range of applications in engineering and the sciences, particularly those involving various integral methods for partial differential equations. Mesh-based methods require a curved mesh for high-order convergence, which can be difficult to reliably obtain on many surfaces, and most meshfree methods require the ability to integrate a set of functions (such as radial basis functions) exactly on the domain of interest; these integrals are generally not known in closed form on most surfaces. We describe two methods for integrating on arbitrary, piecewise-smooth surfaces with or without boundary. Our approaches do not require a particular arrangement of points or an initial triangulation of the surface, making them completely meshfree. We also show how the methods can be extended to handle singular integrals while maintaining high accuracy without changing the point density near singularities.
keywords:
numerical integration, meshfree methods, radial basis functions, surface integration
{AMS}
65D30, 65D32, 65D12, 65N35
1 Introduction
A classical problem in a wide range of scientific disciplines is estimating $\int_{S}f$ from data obtained in a domain $S$. There are a variety of approaches, including solving moment fitting equations [28, 29], interpolating data with radial basis functions (RBFs) [24, 8], integrating polynomials on polyhedra (often for use with finite elements) [11, 18], and Green’s theorem-based methods for dimension reduction [9]. The majority of these methods (except for some RBF methods) typically determine quadrature points and weights for some predetermined class of domains such as polyhedra or domains bounded by algebraic curves. In cases where $f$ is not known everywhere, evaluating $f$ at quadrature points requires another interpolation step. Most methods (including many RBF methods) also require knowledge of the moments of a class of functions on the domain of interest. These moments are generally unknown, except on the simplest domains.
Surfaces present a particular challenge for integration as they are often more difficult to mesh. For meshed methods, it should be noted that simply increasing the degree of the polynomial used to interpolate a function on a standard surface mesh with flat, polygonal elements will not improve the order of convergence beyond second order; the surface itself must also be approximated. This can be done with curved elements, such as those used in high-order surface finite element methods [5], or by using some other approximation of the surface, such as a least-squares approach combined with standard triangular elements as in [21]. We note that actually producing a high-order surface mesh can be challenging and time-consuming, and often involves interpolating a low-order mesh, which may require using a meshfree method (see [36], for example).
Another possible solution is to use a standard quadrature method in the embedding space and discretize surface integrals using a regularization of the Dirac delta. Such approaches can be used with level-set methods (see [6], for example) or when a closest-point representation is available [15], and can be made high order with a suitable choice of regularization [32]. However, these techniques require increasingly expensive discretizations of a neighbourhood of the surface in the embedding space as the accuracy of the regularization is improved. Furthermore, an approximation of the signed distance function for the surface is needed, which may involve additional computation. These methods are also generally only suitable for differentiable closed surfaces for which a level set description is known, excluding surfaces with boundary. Moreover, for surfaces that are only piecewise-smooth, convergence is at a reduced rate.
For meshfree methods, techniques that rely on knowing the exact integral of a class of functions (such as RBFs) on the domain are typically not available, since these exact integrals are often unknown. Meshfree integration on surface point clouds is therefore often limited to Monte Carlo methods, which have been employed for eigenfunction estimation using the weak form of differential operators [10, 35]. Even though the underlying meshfree interpolation technique used is high-order, Monte Carlo integration is not, which limits the overall rate of convergence. Monte Carlo integration also requires knowledge of the distribution from which a point cloud is sampled, which may not be available, and it may not be straightforward to sample points from a surface uniformly. If the points satisfy a minimum distance condition, so that no two points are too close together, a potential integration technique for non-uniform point clouds is to use local projections to the tangent plane and a triangulation to compute Voronoi cell areas (see [16]), though this has a limited order of convergence. Existing high-order meshfree integration methods on surfaces are often restricted to specific surfaces where exact integrals of RBFs are known, such as the sphere [13]. Reeger, Fornberg, and Watts [22] have also introduced a high-order ($\mathcal{O}\left(h^{7}\right)$, where $h$ is the point spacing), RBF-based method for general closed surfaces that requires first obtaining a triangulation of the point cloud. This is essentially a hybrid approach between a meshfree and meshed method in that while a triangulation of the point cloud is required, a meshfree method is ultimately used to compute integrals on each surface patch projected to a triangle with high accuracy. A version for surfaces with boundary has also been developed [23].
Our contribution consists of two super-algebraically convergent methods based on the divergence theorem and norm-minimizing Hermite-Birkhoff interpolation schemes. Such interpolation schemes include Hermite RBFs [26, 7] and minimum norm methods using Fourier series [4, 31]. We review these interpolation methods in Section 2, including a generalized convergence statement in Proposition 2.1, a discussion on a general approach to high-order convergence in 2.2, and two examples in 2.3. We then introduce our integration methods in Section 3. Our approaches do not require specific quadrature points (the points may be randomly scattered and non-uniformly selected) and only require the knowledge of basic domain information such as normal vectors at scattered points on the surface. One method can be used to directly evaluate the ratio of two integrals on a closed surface or a surface with boundary; this is useful when a constant multiple of the integral is all that is needed. For example, the method could be used when the average value of a function on the surface is desired. The second method requires surfaces to have boundary, but we discuss simple strategies for dividing closed surfaces into multiple surfaces with boundary in Subsection 3.5. Unlike the first method, the second method estimates integrals themselves rather than a ratio, which can be used to estimate surface area with high precision. We show that these schemes converge at the same rate as the underlying interpolation methods, regardless of the point distribution; this leads to arbitrarily high-order schemes. Then, we test the integration techniques for a variety of problems on flat domains and on surfaces in Section 4. We conclude by introducing a more general view of norm-minimizing meshfree methods suitable for handling singularities, and then adapt our methods for singular integrals in Subsection 4.2; such integrals appear in boundary integral solutions to PDEs. Code for running the numerical tests in this paper is available at https://github.com/venndaniel/meshfree_integration.
2 Background
We start by giving a brief introduction to solving partial differential equations (PDEs) with meshfree methods. In particular, we look at symmetric, strong-form collocation approaches that choose a Hermite-Birkhoff interpolant that minimizes a norm in a Hilbert space. These approaches can be viewed as constrained optimization problems, which is a vital point for understanding Method 1 in Subsection 3.1.
2.1 Norm-Minimizing Hermite-Birkhoff Interpolants
To solve a PDE, one possible approach is strong-form collocation. In such methods, we search for an approximate solution $\tilde{u}$ that satisfies the PDE on a discrete set of points $\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\tilde{N}}$ contained in a domain $S\subset\mathbb{R}^{m}$. More precisely, if we consider the problem
(1) $\mathcal{F}u$ $=f\text{ on }S,$
where $\mathcal{F}$ is a linear differential operator on $S$ and $f:S\to\mathbb{C}$ is a function, we would search for an approximate solution $\tilde{u}$ such that
(2) $\left(\mathcal{F}\tilde{u}\right)\left(\boldsymbol{x}_{j}\right)=f\left(\boldsymbol{x}_{j}\right)\text{ for each }j\in\left\{1,2,\ldots,\tilde{N}\right\}$.
Boundary conditions can be handled in a number of ways. A standard approach would be to select $\tilde{u}$ from a finite-dimensional space of functions that match the boundary condition of interest. For example, a cosine series could be used for a 1D problem with homogeneous Neumann boundary conditions. However, we are interested in irregular domains and surfaces, where suitable function spaces are generally not known. One possible approach is to simply include the boundary conditions as additional interpolation conditions. That is, if we have the boundary condition
(3) $\mathcal{G}u\biggr|_{\partial S}=g,$
we further require that in addition to satisfying Eq. (2), our approximate solution $\tilde{u}$ satisfies
(4) $\left(\mathcal{G}\tilde{u}\right)\left(\boldsymbol{y}_{j}\right)=g\left(\boldsymbol{y}_{j}\right)\text{ for each }j\in\left\{1,2,\ldots,\tilde{N}_{\partial}\right\}$,
where $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\tilde{N}_{\partial}}\subset\partial S$ is a set of points on the boundary of $S$. The goal is then to ensure that $\mathcal{F}\tilde{u}\to f$ on $S$ and $\mathcal{G}\tilde{u}\to g$ on $\partial S$ in some sense as $\tilde{N},\tilde{N}_{\partial}\to\infty$. If the PDE given by (1) and (3) is suitably well-posed, we may then be able to show that $\tilde{u}\to u$, with the manner of convergence ($L^{2},L^{\infty},H^{p}$, pointwise, etc.) depending on the PDE and method chosen (see, for example, Proposition 11 of [31]).
Of course, simply imposing $\left(\mathcal{F}\tilde{u}\right)\left(\boldsymbol{x}_{j}\right)=f\left(\boldsymbol{x}_{j}\right)$ on each $\boldsymbol{x}_{j}$ is not enough to ensure that $\mathcal{F}\tilde{u}\to f$ as more collocation points are added. A classic counterexample is the interpolation of Runge’s function ($f\left(x\right)=1/\left(1+25x^{2}\right)$) with polynomials of increasing degree on $\left[-1,1\right]$ and evenly spaced points $\left\{\boldsymbol{x}_{j}\right\}$; this corresponds to an identity operator $\mathcal{F}$ and results in an interpolant $\tilde{u}$ that is unbounded as $\tilde{N}\to\infty$. While basic, this counterexample demonstrates one possible manner in which interpolation can fail: the interpolant can become unbounded. This motivates a constrained approach. For some Hilbert space of functions $\mathcal{H}$, we solve
(5) $\text{minimize, over }\tilde{v}\in\mathcal{H}$: $\left\|\tilde{v}\right\|_{\mathcal{H}}$,
subject to $\left(\mathcal{F}\tilde{v}\right)\left(\boldsymbol{x}_{j}\right)=f\left(\boldsymbol{x}_{j}\right)\text{ for each $j$$\in\left\{1,2,\ldots,\tilde{N}\right\},$}$
$\left(\mathcal{G}\tilde{v}\right)\left(\boldsymbol{y}_{j}\right)=g\left(\boldsymbol{y}_{j}\right)\text{ for each $j$$\in\left\{1,2,\ldots,\tilde{N}_{\partial}\right\}.$}$
It is convenient to define a linear operator $\mathcal{L}:\mathcal{H}\to\mathbb{C}^{\tilde{N}+\tilde{N}_{\partial}}$ and some $\boldsymbol{f}\in\mathbb{C}^{\tilde{N}+\tilde{N}_{\partial}}$ such that this problem becomes
(6) $\text{minimize, over }\tilde{v}\in\mathcal{H}$: $\left\|\tilde{v}\right\|_{\mathcal{H}}$,
subject to $\mathcal{L}\tilde{v}=\boldsymbol{f}.$
For this problem to be uniquely solvable, we require $\mathcal{L}$ to be bounded in the $\mathcal{H}\to\mathbb{C}^{\tilde{N}+\tilde{N}_{\partial}}$ sense, which means that both $\mathcal{F}$ and $\mathcal{G}$ evaluated at individual points must be bounded as operators from $\mathcal{H}$ to $\mathbb{C}$. We also need the constraint set to be non-empty, which is often the case due to the finite number of collocation points even if the PDE given by (1) and (3) is not solvable globally. In this case, problem (6) has a few important properties. Let $\tilde{u}\in\mathcal{H}$ be the solution to (6), then $\tilde{u}\in\mathcal{R}\left(\mathcal{L}^{*}\right)=\left(\mathcal{N}\left(\mathcal{L}\right)\right)^{\perp}$ (noting that $\mathcal{R}\left(\mathcal{L}^{*}\right)$ is finite-dimensional and therefore closed). We can therefore find $\tilde{u}$ by solving
(7) $\mathcal{L}\mathcal{L}^{*}\boldsymbol{\beta}$ $=\boldsymbol{f},\text{$\quad$}\tilde{u}=\mathcal{L}^{*}\boldsymbol{\beta}.$
The linear problem given by (7) is finite-dimensional, and $\mathcal{L}\mathcal{L}^{*}$ is symmetric and positive-definite as long as the constraint set is non-empty; $\mathcal{L}\mathcal{L}^{*}$ corresponds to the interpolation matrix $\boldsymbol{\Phi}$ for (Hermite) RBF methods. This gives us the term “symmetric meshfree methods” to refer to methods that arise from the solution of (6), since $\mathcal{L}\mathcal{L}^{*}$ is symmetric even when the operator $\mathcal{F}$ is not. Note that there are also non-symmetric meshfree methods, such as Kansa’s [12], for which the methods and results of this paper do not apply.
Now, as long as $\tilde{u}$ remains bounded as $\tilde{N},\tilde{N}_{\partial}\to\infty$, $\tilde{u}$ converges in $\mathcal{H}$. This is due to the following result in functional analysis. We stated versions of this result specific to meshfree methods with affine constraints in our previous works [31, 30]. The version presented here generalizes these results to convex constraint sets in Hilbert spaces.
Proposition 2.1.
Let $\left\{V_{n}\right\}_{n=1}^{\infty}$ be a collection of closed, non-empty, convex sets in a Hilbert space $\mathcal{H}$ such that $V_{1}\supseteq V_{2}\supseteq\ldots$ and define
$\tilde{u}_{n}=\text{argmin}\left\{\left\|\tilde{u}\right\|_{\mathcal{H}}:\tilde{u}\in V_{n}\right\}.$
Assume $\left\{\left\|\tilde{u}_{n}\right\|_{\mathcal{H}}\right\}_{n=1}^{\infty}$ is bounded by a constant $B>0$, then $\tilde{u}_{n}$ converges in $\mathcal{H}$ to some $\tilde{u}_{\infty}\in\bigcap_{n=1}^{\infty}V_{n}$.
Proof 2.2.
Note that $u_{n}$ is well-defined due to a standard result in functional analysis (see Theorem 3.3-1 of Kreyszig’s text [14], for example), since $V_{n}$ is non-empty, closed, and convex by assumption. Furthermore, $\left\{\left\|\tilde{u}_{n}\right\|_{\mathcal{H}}\right\}_{n=1}^{\infty}$ is non-decreasing, since $V_{n_{1}}\supseteq V_{n_{2}}$ whenever $n_{1}<n_{2}$. For any convex combination $\left(1-\alpha\right)\tilde{u}_{n_{1}}+\alpha\tilde{u}_{n_{2}}$ such that $n_{1}<n_{2}$ and $\alpha\in\left(0,1\right)$,
$\left\|u_{n_{1}}\right\|_{\mathcal{H}}^{2}$ $\leq\left\|\alpha\left(\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}}\right)+\tilde{u}_{n_{1}}\right\|_{\mathcal{H}}^{2}$
$=\left\|u_{n_{1}}\right\|_{\mathcal{H}}^{2}+\alpha^{2}\left\|\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}}\right\|_{\mathcal{H}}^{2}+2\alpha\,\text{Re}\left(\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}},\tilde{u}_{n_{1}}\right)_{\mathcal{H}}$
(8) $\implies 0$ $\leq\alpha\left\|\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}}\right\|_{\mathcal{H}}^{2}+2\,\text{Re}\left(\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}},\tilde{u}_{n_{1}}\right)_{\mathcal{H}}.$
Note that the first line above is because $\tilde{u}_{n_{2}}\in V_{n_{2}}\subseteq V_{n_{1}}$, $\alpha\tilde{u}_{n_{1}}+\left(1-\alpha\right)\tilde{u}_{n_{2}}\in V_{n_{1}}$ by convexity, and $\tilde{u}_{n_{1}}=\text{argmin}\left\{\left\|\tilde{u}\right\|_{\mathcal{H}}:\tilde{u}\in V_{n_{1}}\right\}$. Now, since (8) holds for each $\alpha\in\left(0,1\right)$,
(9) $\text{Re}\left(\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}},\tilde{u}_{n_{1}}\right)_{\mathcal{H}}$ $\geq 0$
(10) $\implies\text{Re}\left(\tilde{u}_{n_{2}},\tilde{u}_{n_{1}}\right)_{\mathcal{H}}$ $\geq\left\|\tilde{u}_{n_{1}}\right\|_{\mathcal{H}}^{2}$
Then,
$\left\|\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}}\right\|_{\mathcal{H}}^{2}$ $=\text{Re}\left(\tilde{u}_{n_{2}},\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}}\right)-\text{Re}\left(\tilde{u}_{n_{1}},\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}}\right)$
[…]
(11) $\leq 2B\left(\left\|\tilde{u}_{n_{2}}\right\|_{\mathcal{H}}-\left\|\tilde{u}_{n_{1}}\right\|_{\mathcal{H}}\right)\text{, since $\left\|\tilde{u}_{n}\right\|_{\mathcal{H}}\leq B$ for each $n$}.$
Finally, $\left\{\left\|\tilde{u}_{n}\right\|_{\mathcal{H}}\right\}_{n=1}^{\infty}$ is bounded and non-decreasing, and therefore converges, which implies $\left\|\tilde{u}_{n_{2}}-\tilde{u}_{n_{1}}\right\|_{\mathcal{H}}\to 0$ as $n_{1},n_{2}\to\infty$ due to (11). The sequence $\left\{\tilde{u}_{n}\right\}_{n=1}^{\infty}$ is therefore Cauchy and must converge to some $\tilde{u}_{\infty}\in\mathcal{H}$, since $\mathcal{H}$ is a Hilbert space. Furthermore, for each $m$, $\left\{\tilde{u}_{n}\right\}_{n=m}^{\infty}\subset V_{m}$, which is closed, so $\tilde{u}_{\infty}\in V_{m}$ as well. Finally, $\tilde{u}_{\infty}\in\bigcap_{n=1}^{\infty}V_{n}$.
We note that the constraint set for problem (6) is closed as long as $\mathcal{L}$ is bounded, which allows us to apply Proposition 2.1. Also note the following corollary of Proposition 2.1.
Corollary 2.3.
Where $\tilde{u}_{n}$ and $\tilde{u}_{\infty}$ are defined as in Proposition 2.1 and the assumptions of Proposition 2.1 hold,
$\tilde{u}_{\infty}=\text{argmin}\left\{\left\|\tilde{u}\right\|_{\mathcal{H}}:\tilde{u}\in\bigcap_{n=1}^{\infty}V_{n}\right\}.$
Proof 2.4.
Let $\tilde{u}_{\text{min}}=\text{argmin}\left\{\left\|\tilde{u}\right\|_{\mathcal{H}}:\tilde{u}\in\bigcap_{n=1}^{\infty}V_{n}\right\}$; note $\bigcap_{n=1}^{\infty}V_{n}$ is non-empty since we already know $\tilde{u}_{\infty}$ exists and is in $\bigcap_{n=1}^{\infty}V_{n}$. Then, $\left\|\tilde{u}_{n}\right\|_{\mathcal{H}}\to\left\|\tilde{u}_{\infty}\right\|_{\mathcal{H}}$ and $\left\|\tilde{u}_{\infty}\right\|_{\mathcal{H}}\geq\left\|\tilde{u}_{\text{min}}\right\|_{\mathcal{H}}\geq\left\|\tilde{u}_{n}\right\|_{\mathcal{H}}$ for each $n$, so $\left\|\tilde{u}_{\infty}\right\|_{\mathcal{H}}=\left\|\tilde{u}_{\text{min}}\right\|_{\mathcal{H}}$. Finally, $\tilde{u}_{\infty}=\tilde{u}_{\text{min}}$ because $\bigcap_{n=1}^{\infty}V_{n}$ is convex and therefore its minimum-norm element is unique.
For our optimization problem applied to PDEs, assume that $\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\infty}$ is dense in $S$ and that $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\infty}$ is dense in $\partial S$. Then, Corollary 2.3 implies that if our solution $\tilde{u}_{\tilde{N},\tilde{N}_{\partial}}$ to the optimization problem (6) remains bounded in $\mathcal{H}$ as $\tilde{N},\tilde{N}_{\partial}\to\infty$, there is at least one solution to the PDE on a dense set of points in $S$ and $\partial S$. If $\mathcal{F}u,\mathcal{G}u$ are continuous for functions $u\in\mathcal{H}$, there is a solution to the PDE on all of $S$ and $\partial S$, and our solutions to (6) converge to $\tilde{u}_{\infty}$ where
$\tilde{u}_{\infty}=\text{argmin}\left\{\left\|\tilde{u}\right\|_{\mathcal{H}}:\mathcal{F}\tilde{u}\biggr|_{S}=f\text{ and }\mathcal{G}\tilde{u}\biggr|_{\partial S}=g\right\}.$
That is, $\tilde{u}_{\tilde{N},\tilde{N}_{\partial}}$ converges to the norm-minimizing solution to the PDE given by (1) and (3). Also note that if at least one strong-form solution to the PDE of (1) and (3) exists and can be extended to a function $u\in\mathcal{H}$, then $u$ is feasible for each (6) and $\left\|\tilde{u}_{\tilde{N},\tilde{N}_{\partial}}\right\|_{\mathcal{H}}\leq\left\|u\right\|_{\mathcal{H}}$ for each $n$, so Proposition 2.1 and Corollary 2.3 apply; $\tilde{u}_{\tilde{N},\tilde{N}_{\partial}}\to\tilde{u}_{\infty}$, where $\tilde{u}_{\infty}$ also solves the PDE of (1) and (3).
2.2 High-Order Bounds on Functions with Scattered Zeros
Before discussing convergence rates, we must recall a couple of definitions.
Definition 2.5.
The fill distance $h_{\text{max}}$ of a set of points $\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\tilde{N}}$ in a domain $S$ is given by
$h_{\text{max}}=\sup_{\boldsymbol{x}\in S}\min_{j\in\left\{1,2,\ldots,\tilde{N}\right\}}\left\|\boldsymbol{x}-\boldsymbol{x}_{j}\right\|_{2}.$
A beneficial aspect of meshfree methods is that they typically converge in a high-order manner with respect to the fill distance. Various RBF interpolation schemes are known to converge super-algebraically when used to approximate functions both in and out of the native Hilbert space $\mathcal{H}$ on certain types of domains (see, for example, Subsection 3.2 of [19] for functions outside the native space). A common requirement is that the domain satisfies an interior cone condition, which we now define.
Definition 2.6.
(See Def. 3.6 from [33]) A domain $U\subset\mathbb{R}^{m}$ satisfies an interior cone condition if there exists an angle $\theta\in\left(0,\frac{\pi}{2}\right)$ and a radius $r>0$ such that for all $\boldsymbol{x}\in U$, there exists a unit vector $\boldsymbol{\xi}\left(\boldsymbol{x}\right)$ such that
$\left\{\boldsymbol{x}+\lambda\boldsymbol{y}:\boldsymbol{y}\in\mathbb{R}^{m},\left\|\boldsymbol{y}\right\|_{2}=1,\boldsymbol{y}\cdot\boldsymbol{\xi}\left(\boldsymbol{x}\right)\geq\cos\theta,\lambda\in\left[0,r\right]\right\}\subset U.$
Bounded Lipschitz domains are an example of domains that satisfy an interior cone condition. One path towards efficiently proving convergence of certain meshfree interpolation schemes is to note that functions with scattered zeros satisfy high-order bounds. This can be proven directly, as in [19], or using existing results from other interpolation schemes, which is the path we have previously taken in [31] using results from [33], summarized in the next result.
Theorem 2.7.
(Adapted from Corollary 6 of [31]) Let $f\in H^{\frac{m}{2}+q+\frac{1}{2}}\left(U\right)$ where $U\subset\mathbb{R}^{m}$ is bounded and satisfies an interior cone condition, and $m\geq 3$ if $q=0$. Assume $f$ is exactly zero on $\left\{x_{k}\right\}_{k=1}^{\tilde{N}}$ with fill distance $h_{\text{max}}$ on $U$, then for all $p\leq q$, there exist constants $C_{m,p},h_{0,m,p}>0$ such that as long as $h_{\text{max}}<h_{0,m,p}$,
$\left\|f\right\|_{L^{\infty}\left(U\right)}\leq C_{m,p}h_{\text{max}}^{p+\frac{1}{2}}\left\|f\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(U\right)}.$
On compact manifolds $S$ with dimension $m$, the same theorem holds, with the sufficient conditions that $S$ can be parametrized by a finite atlas $\left\{\sigma_{k}\right\}_{k=1}^{M}$ with $0<M<\infty$ such that $\sigma_{k}:U_{k}\to S$ is $C^{\left\lceil\frac{m}{2}+q+\frac{1}{2}\right\rceil}$ with bounded derivatives and has a bounded inverse tensor, and each $U_{k}$ is bounded and satisfies an interior cone condition. The theorem may then be applied to $f\circ\sigma_{k}$ on $U_{k}$. The utility of Theorem 2.7 is that if our Hilbert space $\mathcal{H}$ is such that $\left\|\mathcal{F}v\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(S\right)}\leq A\left\|v\right\|_{\mathcal{H}}$ and $\left\|\mathcal{G}v\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(\partial S\right)}\leq B\left\|v\right\|_{\mathcal{H}}$ for each $v\in\mathcal{H}$ and some constants $A,B$, then we can immediately state estimates for $\left\|\mathcal{F}\tilde{u}-f\right\|_{L^{\infty}\left(S\right)}$ and $\left\|\mathcal{G}\tilde{u}-g\right\|_{L^{\infty}\left(\partial S\right)}$ where $\tilde{u}$ is the solution to (5). In particular, if a solution $u$ to the PDE given by (1) and (3) exists and $\left\|\mathcal{F}v\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(S\right)}\leq A\left\|v\right\|_{\mathcal{H}}$ for each $v\in\mathcal{H}$,
$\left\|\mathcal{F}\tilde{u}-f\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(S\right)}$ $=\left\|\mathcal{F}\left(\tilde{u}-u\right)\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(S\right)}\leq A\left\|\tilde{u}-u\right\|_{\mathcal{H}}.$
Then, $\left\|\tilde{u}-u\right\|_{\mathcal{H}}\to 0$ by Proposition 2.1, so if we invoke Theorem 2.7 (assuming $S$ satisfies sufficient conditions), noting that $\mathcal{F}\tilde{u}-f$ has scattered zeros, we have, for $p\leq q$,
$\left\|\mathcal{F}\tilde{u}-f\right\|_{L^{\infty}\left(S\right)}=o\left(h_{\text{max}}^{p+\frac{1}{2}}\right).$
There is then a similar statement for $\left\|\mathcal{G}\tilde{u}-g\right\|_{L^{\infty}\left(\partial S\right)}$. Whether or not a solution $u\in\mathcal{H}$ to the PDE (1) and (3) exists, Theorem 2.7 still implies
$\left\|\mathcal{F}\tilde{u}-f\right\|_{L^{\infty}\left(S\right)}$ $\leq C_{m,p}h_{\text{max}}^{p+\frac{1}{2}}\left\|\mathcal{F}\tilde{u}-f\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(S\right)}$
$\leq C_{m,p}h_{\text{max}}^{p+\frac{1}{2}}\left(A\left\|\tilde{u}\right\|_{\mathcal{H}}+\left\|f\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(S\right)}\right),$
with a similar expression for $\mathcal{G}\tilde{u}-g$ again. Leaving $f,g$, and $S$ fixed (or $f,g$ varying, but bounded), but varying $h_{\text{max}}$, $\left\|\tilde{u}\right\|_{\mathcal{H}}$ either converges to a non-zero value, assuming $f$ or $g$ are non-zero, or diverges to infinity due to Proposition 2.1. With this in mind, we will often assume for simplicity that the method given by (5) is $p^{\text{th}}$-order in the sense that,
(12) $\left\|\mathcal{F}\tilde{u}-f\right\|_{L^{\infty}\left(S\right)}$ $\leq A_{p}h_{\text{max}}^{p}\left\|\tilde{u}\right\|_{\mathcal{H}}$,
$\left\|\mathcal{G}\tilde{u}-g\right\|_{L^{\infty}\left({\partial S}\right)}$ $\leq B_{p}h_{\text{max}}^{p}\left\|\tilde{u}\right\|_{\mathcal{H}}$.
for some constants $A_{p},B_{p}>0$. We again note that (12) holds regardless of whether the PDE given by (1) and (3) is even solvable in $\mathcal{H}$; if it is, then $\left\|\tilde{u}\right\|_{\mathcal{H}}$ is bounded and $\left\|\mathcal{F}\tilde{u}-f\right\|_{L^{\infty}\left(S\right)},$ $\left\|\mathcal{G}\tilde{u}-g\right\|_{L^{\infty}\left({\partial S}\right)}\to 0$. If it is not solvable, $\left\|\tilde{u}\right\|_{\mathcal{H}}$ will diverge, potentially faster than $h_{\text{max}}^{-p}$, so $\left\|\mathcal{F}\tilde{u}-f\right\|_{L^{\infty}\left(S\right)}$ and $\left\|\mathcal{G}\tilde{u}-g\right\|_{L^{\infty}\left({\partial S}\right)}$ could diverge.
2.3 Hilbert Spaces for Meshfree Methods
Recall that $\mathcal{L}$ must be bounded as an operator from $\mathcal{H}$ to $\mathbb{C}^{\tilde{N}+\tilde{N}_{\partial}}$ to guarantee convergence via Proposition 2.1. A straightforward approach to construct a suitable Hilbert space is to use Fourier transforms or series.
Example 2.8.
Consider the Hilbert space of functions on $\mathbb{R}$ given by
$\mathcal{H}:=\left\{u:u\left(x\right)=\int_{-\infty}^{\infty}e^{-\left|\omega\right|+i\omega x}a\left(\omega\right)\,\text{d}\omega,a\in L^{2}\left(\mathbb{R}\right)\right\},$
such that if $u\left(x\right)=\int_{-\infty}^{\infty}e^{-\left|\omega\right|+i\omega x}a\left(\omega\right)\,\text{d}\omega$ and $v\left(x\right)=\int_{-\infty}^{\infty}e^{-\left|\omega\right|+i\omega x}b\left(\omega\right)\,\text{d}\omega$, then the inner product on $\mathcal{H}$ is given by $\left(u,v\right)_{\mathcal{H}}:=\left(a,b\right)_{L^{2}\left(\mathbb{R}\right)}.$ Now consider the interpolation problem on $\left\{x_{j}\right\}_{j=1}^{\tilde{N}}\subset S\subset\mathbb{R}$.
(13) $\text{minimize, over }u\in\mathcal{H}:$ $\left\|u\right\|_{\mathcal{H}},$
subject to $u\left(x_{j}\right)=f_{j}\text{ for each $j$$\in\left\{1,2,\ldots,\tilde{N}\right\},$}$
for some $\boldsymbol{f}\in\mathbb{C}^{\tilde{N}}$. Define $\mathcal{L}:\mathcal{H}\to\mathbb{C}^{\tilde{N}}$ by $\left(\mathcal{L}u\right)_{j}:=u\left(x_{j}\right)$. Then,
$\left(\mathcal{L}^{*}\boldsymbol{\beta}\right)\left(x\right)$ $=\sum_{j=1}^{\tilde{N}}\beta_{j}\int_{-\infty}^{\infty}e^{-2\left|\omega\right|+i\omega\left(x-x_{j}\right)}\,\text{d}\omega=\sum_{j=1}^{\tilde{N}}\beta_{j}\frac{4}{4+\left(x-x_{j}\right)^{2}}.$
That is, $\mathcal{L}^{*}\hat{\boldsymbol{e}}_{j}$ corresponds to the RBF $\phi$ such that $\phi\left(x-x_{j}\right):=4/\left(4+\left(x-x_{j}\right)^{2}\right)$: an inverse quadratic RBF. Indeed, the RBF interpolation matrix $\boldsymbol{\Phi}$ for this problem is given by $\Phi_{jk}=4/\left(4+\left(x_{j}-x_{k}\right)^{2}\right)=\left(\mathcal{L}\mathcal{L}^{*}\right)_{jk}.$
Example 2.9.
Consider placing a domain $S$ inside a box $\Omega\in\mathbb{R}^{m}$. We consider the set of periodic Fourier basis functions on $\Omega$: $\left\{\phi_{n}\right\}_{n=1}^{\infty}$ where $\phi_{n}\left(\boldsymbol{x}\right)=e^{i\boldsymbol{\omega}_{n}\cdot\boldsymbol{x}}$ for frequencies $\boldsymbol{\omega}_{n}$ such that $\phi_{n}$ is periodic on $\Omega$. Then, define a Hilbert space $\mathcal{H}$ by
(14) $\mathcal{H}:=\left\{\sum_{n=1}^{\infty}a_{n}d_{n}^{-\frac{1}{2}}\phi_{n}:a\in\ell^{2}\right\},$
where $d=\left\{d_{n}\right\}_{n=1}^{\infty}>0$ is a series that must be chosen. In the case that $\mathcal{F}$ and $\mathcal{G}$ from Subsection 2.1 are linear differential operators of at most order $p$ with bounded coefficient functions, the condition for $\mathcal{L}$ to be bounded (and therefore the method to converge) is that $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$, where $\left\|\boldsymbol{\omega}\right\|_{2}:=\left\{\left\|\boldsymbol{\omega}_{n}\right\|_{2}\right\}_{n=1}^{\infty}$ is a sequence of Euclidean norms of the Fourier frequencies (see [31] for further discussion). Another possible construction of this Hilbert space is
$\mathcal{H}:=\left\{\sum_{n=1}^{\infty}\hat{u}_{n}\phi_{n}:\sum_{n=1}^{\infty}d_{n}\left|\hat{u}_{n}\right|^{2}<\infty\right\}.$
In this construction, it is clear that choosing $d_{n}=\left(1+\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{2}\right)^{p}$ corresponds to minimum Sobolev norm interpolation (see [4], for example). The inner product is
$\left(\sum_{n=1}^{\infty}a_{n}d_{n}^{-\frac{1}{2}}\phi_{n},\sum_{n=1}^{\infty}b_{n}d_{n}^{-\frac{1}{2}}\phi_{n}\right)_{\mathcal{H}}:=\left(a,b\right)_{\ell^{2}}.$
In dimensions greater than one, it is useful to make $d$ separable; for example, in 2D, we could choose $d_{n}:=d_{x,n_{x}}d_{y,n_{y}}$, where $\boldsymbol{\omega}_{n}=\omega_{x,n_{x}}\hat{\boldsymbol{x}}+\omega_{y,n_{y}}\hat{\boldsymbol{y}}$. Then, in the case $\mathcal{F},\mathcal{G}$ are identities (function interpolation), it can be shown that
(15) $\mathcal{L}^{*}\hat{\boldsymbol{e}}_{j}$ $=\sum_{n=1}^{\infty}d_{n}^{-1}e^{i\boldsymbol{\omega}_{n}\cdot\left(\boldsymbol{x}-\boldsymbol{x}_{j}\right)}=\left(\sum_{n_{x}=1}^{\infty}d_{x,n_{x}}^{-1}e^{i\omega_{x,n_{x}}\left(x-x_{j}\right)}\right)\left(\sum_{n_{y}=1}^{\infty}d_{y,n_{y}}^{-1}e^{i\omega_{y,n_{y}}\left(y-y_{j}\right)}\right).$
This is useful because, in practice, we will always be using a truncated Fourier series. When $d_{n}$ is separable, we can evaluate $\left(\mathcal{L}^{*}\hat{\boldsymbol{e}}_{j}\right)\left(x,y\right)$ with $N_{\omega}=N_{\omega,x}N_{\omega,y}$ Fourier basis functions in $\mathcal{O}\left(N_{\omega,x}+N_{\omega,y}\right)$ operations, rather than $\mathcal{O}\left(N_{\omega,x}N_{\omega,y}\right)$. This means we can typically form the $\boldsymbol{\Phi}=\boldsymbol{VV}^{*}$ matrix in $\mathcal{O}\left(N_{\omega}^{\frac{1}{m}}\left(\tilde{N}+\tilde{N}_{\partial}\right)^{2}\right)$ time for a series in $m$ dimensions rather than the $\mathcal{O}\left(N_{\omega}\left(\tilde{N}+\tilde{N}_{\partial}\right)^{2}\right)$ of computing $\boldsymbol{VV}^{*}$ directly. However, the $\boldsymbol{\Phi}$ matrix is poorly conditioned, so this approach is typically limited to 5 or 6 digits of accuracy.
After truncating the Fourier series, the optimization problem becomes
(16) minimize: $\left\|\boldsymbol{a}\right\|_{2},$
subject to $\boldsymbol{Va}=\boldsymbol{f},$
where $\boldsymbol{V}$ is the matrix such that $\boldsymbol{Va}\in\mathbb{C}^{\tilde{N}+\tilde{N}_{\partial}}$ is a vector consisting of the values of $\mathcal{F}\left(\sum_{n=1}^{N_{\omega}}a_{n}d_{n}^{-\frac{1}{2}}\phi_{n}\right)$ and $\mathcal{G}\left(\sum_{n=1}^{N_{\omega}}a_{n}d_{n}^{-\frac{1}{2}}\phi_{n}\right)$ on $\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\tilde{N}}$ and $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\tilde{N}_{\partial}}$, respectively. In this case, the matrix $\boldsymbol{\Phi}$ for RBF methods corresponds to $\boldsymbol{VV}^{*}$; one could solve $\boldsymbol{VV}^{*}\boldsymbol{\beta}=\boldsymbol{f}$, then the solution to (16) is given by $\boldsymbol{a}=\boldsymbol{V}^{*}\boldsymbol{\beta}$. As stated previously, this approach is poorly conditioned, however, and superior results can generally be obtained from solving (16) using a singular value decomposition or complete orthogonal decomposition. That said, if speed while using a large number of Fourier basis functions is the main concern, forming $\boldsymbol{VV}^{*}=\boldsymbol{\Phi}$ using a separable $d$ as described around Eq. (15) will be less computationally intensive than solving (16) directly. Solving the positive definite system $\boldsymbol{\Phi}\boldsymbol{\beta}=\boldsymbol{f}$ requires $\mathcal{O}\left(\left(\tilde{N}+\tilde{N}_{\partial}\right)^{3}\right)$ operations using a Cholesky decomposition or another direct algorithm, while the cost of forming a decomposition of $\boldsymbol{V}$ is $\mathcal{O}\left(N_{\omega}\left(\tilde{N}+\tilde{N}_{\partial}\right)^{2}\right)$; $N_{\omega}$ may be significantly larger than $\tilde{N}+\tilde{N}_{\partial}$ to ensure $\boldsymbol{V}$ is full rank. See [31] for further discussion on using these Hilbert spaces for solving surface PDEs.
3 Methods and Analysis
We now present two methods for estimating $\int_{S}f$ on irregular flat domains and surfaces. Throughout this section, we assume that $\mathcal{H}$ is a Hilbert space associated with a particular order of convergence for solving a Poisson problem using a constrained optimization approach (5). Specifically, we assume that there exist constants $p,A_{p},B_{p}>0$ such that if $u\in\mathcal{H}$ and $\Delta_{S}u-f$ has scattered zeros on a set of points $\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\tilde{N}}\subset S$ with fill distance $h_{\text{max}}$, then $\left\|\Delta_{S}u-f\right\|_{L^{\infty}\left(S\right)}\leq A_{p}h_{\text{max}}^{p}\left\|u\right\|_{\mathcal{H}}$ for a sufficiently smooth $f$ and small enough $h_{\text{max}}$. Furthermore, if $\hat{\boldsymbol{n}}_{\partial S}\cdot\nabla_{S}u=0$ on a set of scattered points $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\tilde{N}}\subset\partial S$ with fill distance $h_{\text{max}}$, then $\left\|\hat{\boldsymbol{n}}_{\partial S}\cdot\nabla_{S}u\right\|_{L^{\infty}\left(\partial S\right)}\leq B_{p}h_{\text{max}}^{p}\left\|u\right\|_{\mathcal{H}}$. See the discussion leading up to (12) for an explanation of these assumptions.
3.1 Method 1: Poisson Solvability
We rely on Proposition 2.1 to construct our first method. Let $S\subset\Omega$ be the domain of integration, which may be flat or a surface, and let $g$ be a real function so that $\int_{S}g>0$ is known or can be easily estimated (for example, a function with compact support contained in a flat domain $S$). Assume we want to know $\int_{S}f$, where $f$ is a real integrable function on $S$. Then consider the problem
(17) $\Delta_{S}u$ $=f-cg,\text{$\quad$}\hat{\boldsymbol{n}}_{\partial S}\cdot\nabla_{S}u\biggr|_{\partial S}=0,$
where $\hat{\boldsymbol{n}}_{S}$ is the normal vector to $S$ and $c$ is some constant. If $S$ is a closed surface, the boundary condition is simply removed. The problem (17) is only solvable when $\int_{S}\left(f-cg\right)=0$, or equivalently, when
(18) $c=c_{*}:=\frac{\int_{S}f}{\int_{S}g}.$
We can then discretize problem (17). Let $\mathcal{H}$ be a Hilbert space satisfying the assumptions at the beginning of this section. Also assume that $f$ and $g$ are smooth enough so that the solution to (17) for the value of $c$ given by (18) can be extended to some $u_{c_{*}}\in\mathcal{H}$. Let $\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\infty}$ be dense in $S$, and let $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\infty}$ be dense in $\partial S$. Then consider the problem:
(19) $\text{minimize, over }u\in\mathcal{H}:$ $\left\|u\right\|_{\mathcal{H}},$
subject to $\Delta_{S}u\left(\boldsymbol{x}_{j}\right)=f\left(\boldsymbol{x}_{j}\right)-cg\left(\boldsymbol{x}_{j}\right)\text{ for each $j\in\left\{1,2,\ldots,\tilde{N}\right\}$}$,
$\hat{\boldsymbol{n}}_{\partial S}\cdot\nabla_{S}u\left(\boldsymbol{y}_{j}\right)=0\text{, for each }j\in\left\{1,2,\ldots\tilde{N}_{\partial}\right\}$.
See Subsection 3.4 for the specifics of evaluating $\Delta_{S}u$ on surfaces. We set this up so that the minimizing solution is $\tilde{u}=\mathcal{L}^{*}\boldsymbol{\beta}$, where $\mathcal{L}^{*}$ is as in Subsection 2.1 and $\boldsymbol{\Phi}:=\mathcal{L}\mathcal{L}^{*}$. Then, we can write:
[…]
Now, this will only be bounded as $\tilde{N},\tilde{N}_{\partial}\to\infty$ when $\int_{S}\left(f-cg\right)=0$ using Proposition 2.1. Then, as $\tilde{N},\tilde{N}_{\partial}\to\infty$, the minimizer $\left(\boldsymbol{f}-c\boldsymbol{g}\right)^{*}\boldsymbol{\Phi}^{-1}\left(\boldsymbol{f}-c\boldsymbol{g}\right)$ as a function of $c$ must approach the value of $c$ where $\int_{S}\left(f-cg\right)=0$. This minimum (assuming $\boldsymbol{\Phi}$ is real) is at
(20) $\tilde{c}_{\tilde{N}}:=\frac{\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{f}}{\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{g}},$
where we note that $\boldsymbol{\Phi}^{-1}$ is symmetric. Therefore, as $\tilde{N},\tilde{N}_{\partial}\to\infty$,
(21) $\int_{S}g\frac{\boldsymbol{\boldsymbol{g}}^{*}\boldsymbol{\Phi}^{-1}}{\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{g}}\boldsymbol{f}\to\int_{S}f.$
So, we have a linear operator to estimate $\int_{S}f$, with real weights as long as $\boldsymbol{\Phi}^{-1}$ is real (which is typical for a reasonable choice of $\mathcal{H}$). To be specific, let $\boldsymbol{w}^{*}=\int_{S}g\boldsymbol{\boldsymbol{g}}^{*}\boldsymbol{\Phi}^{-1}/\left(\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{g}\right)$, then $\boldsymbol{w}^{*}\boldsymbol{f}\to\int_{S}f$; $\boldsymbol{w}^{*}$ gives integration weights.
We may also want similar expressions to (20) and (21) in terms of inner products in $\mathcal{H}$. Let $v_{g}^{\left(\tilde{N}\right)}$ and $v_{f}^{\left(\tilde{N}\right)}$ be the solutions to
$\text{minimize, over }u\in\mathcal{H}:$ $\left\|u\right\|_{\mathcal{H}}$
subject to $\Delta_{S}u\left(\boldsymbol{x}_{j}\right)=F\left(\boldsymbol{x}_{j}\right)\text{ for each $j\in\left\{1,2,\ldots,\tilde{N}\right\}$}$
$\hat{\boldsymbol{n}}_{\partial S}\cdot\nabla_{S}u\left(\boldsymbol{y}_{j}\right)=0\text{, for each }j\in\left\{1,2,\ldots\tilde{N}_{\partial}\right\}$,
where $F=g$ and $F=f$ to produce $v_{g}^{\left(\tilde{N}\right)}$ and $v_{f}^{\left(\tilde{N}\right)}$, respectively. These may correspond to problems on the domain $S$ that are not actually solvable, but the discretized version will still admit a unique solution. Then, as $\tilde{N},\tilde{N}_{\partial}\to\infty$,
$\tilde{c}_{\tilde{N}}=\frac{\left(v_{g}^{\left(\tilde{N}\right)},v_{f}^{\left(\tilde{N}\right)}\right)_{\mathcal{H}}}{\left\|v_{g}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}^{2}},\quad\int_{S}g\frac{\left(v_{g}^{\left(\tilde{N}\right)},v_{f}^{\left(\tilde{N}\right)}\right)_{\mathcal{H}}}{\left\|v_{g}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}^{2}}\to\int_{S}f.$
3.1.1 Analysis
We now study the convergence rate of this method; in particular, we show that it inherits the convergence rate of the underlying meshfree PDE method. Again assuming that fcgf-cg is sufficiently smooth and $c$ is in any bounded interval, let $u_{c}^{\left(\tilde{N}\right)}$ be the solution to (19) for a value of $c$ and $\tilde{N}$, then
(22) $\left|\Delta_{S}u_{c}^{\left(\tilde{N}\right)}-\left(f-cg\right)\right|$ $\leq A_{p}h_{\text{max}}^{p}\left\|u_{c}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}\text{ on }S,$
[…]
It is important to note again here that $\left\|u_{c}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}$ may diverge as $\tilde{N}\to\infty$ and $h_{\text{max}}\to 0$; in fact it must diverge if (17) is not solvable for some choice of $c$ by Proposition 2.1. Then, where $\left|\partial S\right|:=\int_{\partial S}1$,
(24) $\left|\int_{S}\Delta_{S}u_{c}^{\left(\tilde{N}\right)}\right|=\left|\int_{\partial S}\hat{\boldsymbol{n}}_{\partial S}\cdot\nabla_{S}u\right|\leq\left|\partial S\right|B_{p}h_{\text{max}}^{p}\left\|u_{c}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}$.
Using (22), (23), and (24),
$\left|\int_{S}\left(f-cg\right)\right|$ $\leq\left(\left|S\right|A_{p}+\left|\partial S\right|B_{p}\right)h_{\text{max}}^{p}\left\|u_{c}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}$.
Then, there exists a constant $\tilde{A}_{p}$ so that
$\left|c-\frac{\int_{S}f}{\int_{S}g}\right|\leq\left|\int_{S}g\right|^{-1}\tilde{A}_{p}h_{\text{max}}^{p}\left\|u_{c}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}$.
In the case that $c=\int_{S}f/\int_{S}g$ and the PDE is solvable, this tells us nothing. However, this holds even when the PDE is not solvable, which gives us an estimate for the rate at which $\left\|u^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}$ diverges:
$\left\|u_{c}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}\geq\frac{h_{\text{max}}^{-p}\left|\int_{S}g\right|}{\tilde{A}_{p}}\left|c-\frac{\int_{S}f}{\int_{S}g}\right|.$
Now assume that an exact, extended, strong-form solution $u_{c_{*}}\in\mathcal{H}$ for $c_{*}=\int_{S}f/\int_{S}g$ exists for problem (17). In this case, $\left\|u_{c_{*}}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}\leq\left\|u_{c_{*}}\right\|_{\mathcal{H}}$ since $u_{c_{*}}$ is feasible for (19). Recall that $\tilde{c}_{\tilde{N}}=\left(\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{f}\right)/\left(\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{g}\right)$ is the minimizer of $\left\|u_{c}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}$ as a function of $c$ and that we already know that $\tilde{c}_{\tilde{N}}\to c_{*}$ from Eq. (21). Then,
$\left\|u_{c_{*}}\right\|_{\mathcal{H}}\geq\left\|u_{c_{*}}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}\geq\left\|u_{\tilde{c}_{\tilde{N}}}^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}\geq\frac{h_{\text{max}}^{-p}\left|\int_{S}g\right|}{\tilde{A}_{p}}\left|\tilde{c}_{\tilde{N}}-c_{*}\right|.$
So, we conclude that $\tilde{c}_{\tilde{N}}\to c_{*}$ at a rate $\mathcal{O}\left(h_{\text{max}}^{p}\right)$ as $h_{\text{max}}\to 0$, since
$\left|\tilde{c}_{\tilde{N}}-c_{*}\right|\leq\tilde{A}_{p}\left|\int_{S}g\right|^{-1}h_{\text{max}}^{p}\left\|u_{c_{*}}\right\|_{\mathcal{H}}.$
Noting that $\tilde{c}_{\tilde{N}}\int_{S}g$ is our estimate for $\int_{S}f$ and that $c_{*}=\int_{S}f/\int_{S}g$, we also have
$\left|\tilde{c}_{\tilde{N}}\int_{S}g-\int_{S}f\right|\leq\tilde{A}_{p}h_{\text{max}}^{p}\left\|u_{c_{*}}\right\|_{\mathcal{H}}=\mathcal{O}\left(h_{\text{max}}^{p}\right).$
This estimate depends only on the norm of the PDE solution for $c=c_{*}$ and tells us that convergence of the integral estimate is at the rate at which $\Delta_{S}u_{c_{*}}^{\left(\tilde{N}\right)}$ and $\hat{\boldsymbol{n}}_{\partial S}\cdot\nabla_{S}u$ converge when solving the PDE for $c=c_{*}$.
3.2 Method 2: Dimension Reduction
An alternative approach for estimating $\int_{S}f$ is to consider
(25) $\Delta_{S}u=f,$
where $S$ is a manifold with boundary. If $S$ is without boundary, we divide it into submanifolds with boundary; this is discussed in Subsection 3.5. We now apply the divergence theorem:
(26) $\int_{S}f=\int_{S}\Delta_{S}u=\int_{\partial S}\nabla_{S}u\cdot\hat{\boldsymbol{n}}_{\partial S},$
where $\hat{\boldsymbol{n}}_{\partial S}$ is the conormal vector along $\partial S$. That is, $\hat{\boldsymbol{n}}_{\partial S}$ is tangential to $S$ and normal to $\partial S$. It is useful to note that we could use any boundary condition for (25). For the meshfree methods considered in this paper, we do not require any boundary conditions at all. According to Proposition 2.1 and Corollary 2.3, as long as (25) admits at least one solution in $\mathcal{H}$, a suitable symmetric meshfree method will produce solutions that converge to the norm-minimizing solution to (25) in $\mathcal{H}$. This emphasizes a key feature of these sorts of meshfree methods: problems that are ill-posed by virtue of admitting multiple solutions become well-posed when the norm-minimization condition is imposed.
Once the divergence theorem is applied to obtain (26), we are left with an integral on a lower-dimensional manifold. The process can then be repeated until only a line integral remains. The discretized problem for (25) is, for a point cloud $\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\tilde{N}}\subset S$,
(27) $\text{minimize, over }u\in\mathcal{H}:$ $\left\|u\right\|_{\mathcal{H}}$
subject to $\Delta_{S}u\left(\boldsymbol{x}_{j}\right)=f\left(\boldsymbol{x}_{j}\right)\text{ for each $j\in\left\{1,2,\ldots,\tilde{N}\right\}$}$.
Let $u^{\left(\tilde{N}\right)}$ be the solution to this problem. If $f,S$, and the functions in $\mathcal{H}$ are sufficiently smooth, then according to (12), we have
(28) $\left|\Delta_{S}u^{\left(\tilde{N}\right)}-f\right|$ $\leq A_{p}h_{\text{max}}^{p}\left\|u^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}\text{ on }S,$
for some $p,A_{p}>0$. We immediately have
$\left|\int_{S}f-\int_{S}\Delta_{S}u^{\left(\tilde{N}\right)}\right|=\left|\int_{S}f-\int_{\partial S}\nabla_{S}u^{\left(\tilde{N}\right)}\cdot\hat{\boldsymbol{n}}_{\partial S}\right|$ $\leq A_{p}\left|S\right|h_{\text{max}}^{p}\left\|u^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}$,
where $\left|S\right|:=\int_{S}1$. This can be repeated for each dimension until $\int_{\partial S}\nabla_{S}u^{\left(\tilde{N}\right)}\cdot\hat{\boldsymbol{n}}_{\partial S}$ is simply a line integral.
3.3 Line Integrals
Line integrals can be evaluated by a wide variety of methods. Curves are much simpler to parametrize than surfaces, so a standard quadrature scheme could be implemented. However, we could also continue to use meshfree methods by considering the first-order problem
(29) $\nabla u\cdot\hat{\boldsymbol{t}}_{C}=f\text{ on }C,$
where $C$ is an open curve of interest and $\hat{\boldsymbol{t}}_{C}$ denotes its unit tangent vector. If we have a closed curve, such as the boundary of a surface, we first divide it into two open curves. Then, $\int_{C}f=u\left(\boldsymbol{b}\right)-u\left(\boldsymbol{a}\right)$, where $\boldsymbol{a}$ and $\boldsymbol{b}$ are the endpoints of $C$. To discretize (29), we use
(30) $\text{minimize, over }u\in\mathcal{H}:$ $\left\|u\right\|_{\mathcal{H}}$
subject to $\left(\nabla u\cdot\hat{\boldsymbol{t}}_{C}\right)\left(\boldsymbol{x}_{j}\right)=f\left(\boldsymbol{x}_{j}\right)\text{ for each $j\in\left\{1,2,\ldots,\tilde{N}\right\}$}$,
where $\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\tilde{N}}$ is now a point cloud on $C$. Once again, with suitable smoothness assumptions, it is straightforward to show that
$\left|\int_{C}f-\left(u^{\left(\tilde{N}\right)}\left(\boldsymbol{b}\right)-u^{\left(\tilde{N}\right)}\left(\boldsymbol{a}\right)\right)\right|\leq A_{p}h_{\text{max}}^{p}\left\|u^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}$,
for some $A_{p}>0$, where $u^{\left(\tilde{N}\right)}$ is the solution to (30) and $p$ depends on the smoothness of $f,S$, and the functions in $\mathcal{H}$.
3.4 Solving Surface Poisson Problems
A complication of solving (19) or (27) on surfaces is that it is not immediately obvious how to compute $\Delta_{S}u$ on a surface; $u$ is a function on a domain $\Omega\supset S$, not on the surface itself, and $\Delta u\biggr|_{S}\neq\Delta_{S}u$ in general. To address this, we note (see Lemma 1 of [34])
(31) $\Delta_{S}u=\Delta u-\kappa\hat{\boldsymbol{n}}_{S}\cdot\nabla u-\hat{\boldsymbol{n}}_{S}\cdot\left(D^{2}u\right)\hat{\boldsymbol{n}}_{S}\text{ on }S,$
where $\hat{\boldsymbol{n}}_{S}$ is the unit normal vector to the surface $S$, $D^{2}u$ is the Hessian, and $\kappa=\nabla_{S}\cdot\hat{\boldsymbol{n}}_{S}$ is the sum of principal curvatures, which is often referred to as the mean curvature in surface PDE communities (the usual mean curvature would be $\kappa/2$ for a surface in this notation). Examining (31), we note that knowing or estimating $\hat{\boldsymbol{n}}_{S}$ is unavoidable. For surfaces described solely by point clouds, $\hat{\boldsymbol{n}}_{S}$ must be estimated; this is a common problem in point cloud processing, and can be handled by fitting a local level set to the point cloud. This could be done with a local version of the algorithms described in [3, 17]. A local high-order parametrization could also be computed, such as in [2]. An analytic expression for $\kappa$ is readily available if $S$ is described by a piecewise parametrization or a level set. We also note that the sign of $\hat{\boldsymbol{n}}_{S}$ is not needed as long as the sign of $\kappa\hat{\boldsymbol{n}}_{S}$ is accurate. This is noteworthy since $\kappa\hat{\boldsymbol{n}}_{S}$ can be estimated locally even when the surface is non-orientable or when a consistent orientation is difficult to obtain.
We use either of the two optimization problems below to impose $\Delta_{S}u=f$ on the point cloud. First, for when both $\kappa$ and $\hat{\boldsymbol{n}}_{S}$ are known or can be computed:
(32) $\text{minimize, over }u\in\mathcal{H}:$ $\left\|u\right\|_{\mathcal{H}}$
subject to $\Delta u-\kappa\hat{\boldsymbol{n}}_{S}\cdot\nabla u-\hat{\boldsymbol{n}}_{S}\cdot\left(D^{2}u\right)\hat{\boldsymbol{n}}_{S}=f\text{ on }\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\tilde{N}}$.
This approach is particularly useful for implicit surfaces, where a parametrization may not be known and meshing may be difficult, but analytic expressions for $\kappa$ and $\hat{\boldsymbol{n}}_{S}$ can be obtained. The other approach, for when only $\hat{\boldsymbol{n}}_{S}$ can be determined, is given by:
(33) $\text{minimize, over }u\in\mathcal{H}:$ $\left\|u\right\|_{\mathcal{H}}$
subject to $\Delta u-\hat{\boldsymbol{n}}_{S}\cdot\left(D^{2}u\right)\hat{\boldsymbol{n}}_{S}=f\text{ on }\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\tilde{N}}$,
$\hat{\boldsymbol{n}}_{S}\cdot\nabla u=0\text{ on }\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\tilde{N}}$.
We note that it is unnecessary to impose both $\hat{\boldsymbol{n}}_{S}\cdot\nabla u\left(\boldsymbol{x}_{j}\right)=0$ and $\hat{\boldsymbol{n}}_{S}\cdot\left(D^{2}u\right)\hat{\boldsymbol{n}}_{S}=0$ here, although this has been done before in other surface PDE methods [20] and would work here as well. However, imposing both these conditions results in larger matrices, a longer computation time, and a smaller space of feasible interpolants. A smaller constraint set then leads to a larger value of $\left\|\tilde{u}\right\|_{\mathcal{H}}$ for the interpolant solution, which in turn can result in a less accurate solution. See [31] for more discussion on solving surface PDEs with this approach.
3.5 Creating Surface Subdomains
3.5.1 Voronoi Cells
For the method described in Subsection 3.2, we require $S$ to be with boundary. If $S$ is without boundary, we must write it as the union of subdomains with boundary: $S=\bigcup_{j=1}^{N_{S}}S_{j}$. We would like this process to be completely automated. A very simple process is to divide $S$ into a number of Voronoi cells. Using a sparse point cloud $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{N_{S}}$, we define
$S_{j}:=\left\{\boldsymbol{x}\in S:\left\|\boldsymbol{x}-\boldsymbol{y}_{j}\right\|_{2}\leq\min_{k\neq j}\left\|\boldsymbol{x}-\boldsymbol{y}_{k}\right\|_{2}\right\},$
where $\left\|\cdot\right\|_{2}$ is simply the Euclidean norm in the embedding space (e.g. if $S$ is a surface embedded in $\mathbb{R}^{3}$, then the Voronoi cells use the Euclidean norm in $\mathbb{R}^{3}$, not the metric on $S$). One motivation for using Voronoi cells is that the necessary vectors on $\partial S_{j}$ are straightforward to compute. Specifically, $\hat{\boldsymbol{t}}_{\partial S_{j}}$ is in the intersection of the tangent spaces of $S$ and the plane $P_{jk}$ that separates $S_{j}$ from a neighbouring Voronoi cell $S_{k}$. For surfaces embedded in $\mathbb{R}^{3}$, this means $\hat{\boldsymbol{t}}_{\partial S_{j}}\propto\hat{\boldsymbol{n}}_{S}\times\hat{\boldsymbol{n}}_{P_{jk}}$. The outward normal to $\partial S_{j}$ is then given by $\hat{\boldsymbol{n}}_{\partial S_{j}}\propto\hat{\boldsymbol{t}}_{\partial S_{j}}\times\hat{\boldsymbol{n}}_{S}$.
Point placement in $S_{j}$ can then be automated using algorithms A.1 or A.2 from [30], or any other existing approach for generating point clouds (see [25] for a review). As long as $S_{j}$ is diffeomorphic to a simply connected subset of the plane such that $\partial S_{j}$ consists of piecewise differentiable segments (the boundaries between $S_{j}$ and neighbouring Voronoi cells), points can be placed on $\partial S_{j}$ by placing evenly spaced points between the adjacent corners of $S_{j}$, then moving these points along the plane separating $S_{j}$ from a neighbouring Voronoi cell until they intersect $S$ (via Newton’s method if $S$ is a level set, for example). Integrals on $\partial S_{j}$ are estimated using the method from Subsection 3.3 on each piecewise differentiable segment.
3.5.2 Planar Boundary
A simple approach to divide closed surfaces into two subdomains is to use a single plane defined by $P=\left\{\boldsymbol{x}\in\mathbb{R}^{3}:\boldsymbol{p}\cdot\boldsymbol{x}-c=0\right\}$ for some constant $c$. In this case, the two subdomains are given by
$S_{\pm}=\overline{\left\{\boldsymbol{x}\in S:\text{sign}\left(\boldsymbol{p}\cdot\boldsymbol{x}-c\right)=\pm 1\right\}},$
and the boundaries $\partial S_{\pm}$ are closed planar curves. A complication that can arise for surfaces of non-zero genus is that each boundary may consist of a collection of disjoint curves. We note here that the meshfree approach is primarily advantageous for surfaces where high-order meshing may be challenging; integrals on curves may be computed more simply by other means. A straightforward technique for identifying and integrating over multiple planar, closed curves is as follows:
  1. 1.
    Generate points $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\tilde{N}_{\partial}}$ on $\partial S_{\pm}$. This can be done using algorithm A.1 from [30] modified to constrain points to the plane via projection.
  2. 2.
    For each point in $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\tilde{N}_{\partial}}$ , find its nearest neighbour $\boldsymbol{y}_{j}^{\left(1\right)}\in\left\{\boldsymbol{y}_{k}\right\}_{k=1,k\neq j}^{\tilde{N}_{\partial}}$.
  3. 3.
    For each point in $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\tilde{N}_{\partial}}$, find the nearest point $\boldsymbol{y}_{j}^{\left(2\right)}$ such that $\left(\boldsymbol{y}_{j}-\boldsymbol{y}_{j}^{\left(1\right)}\right)\cdot\left(\boldsymbol{y}_{j}-\boldsymbol{y}_{j}^{\left(2\right)}\right)<0$ (so that $\boldsymbol{y}_{j}^{\left(1\right)}$ and $\boldsymbol{y}_{j}^{\left(2\right)}$ are on opposite sides of $\boldsymbol{y}_{j}$). That is, $\boldsymbol{y}_{j}^{\left(2\right)}:=\text{argmin}\left\{\left\|\boldsymbol{y}-\boldsymbol{y}_{j}\right\|:\boldsymbol{y}\in\left\{\boldsymbol{y}_{k}\right\}_{k=1}^{\tilde{N}_{\partial}},\left(\boldsymbol{y}_{j}-\boldsymbol{y}_{j}^{\left(1\right)}\right)\cdot\left(\boldsymbol{y}_{j}-\boldsymbol{y}\right)<0\right\}$.
  4. 4.
    For each trio $\boldsymbol{y}_{j},\boldsymbol{y}_{j}^{\left(1\right)},\boldsymbol{y}_{j}^{\left(2\right)}$, construct the circle or line passing through each of the three points. Parametrize circles by the angle $\theta$ between the points on the circle, the circle’s centre, and $\boldsymbol{y}_{j}$. Parametrize lines by $s$ where $s\left(\boldsymbol{x}\right)=\left(\boldsymbol{x}-\boldsymbol{y}_{j}\right)\cdot\left(\boldsymbol{y}_{j}^{\left(1\right)}-\boldsymbol{y}_{j}\right)/\left\|\boldsymbol{y}_{j}^{\left(1\right)}-\boldsymbol{y}_{j}\right\|_{2}$ (signed distance from $\boldsymbol{y}_{j}$).
  5. 5.
    Compute weights for each trio so that quadratic functions in $\theta$ or $s$ are integrated exactly on the arc/line between $\boldsymbol{y}_{j}^{\left(1\right)}$ and $\boldsymbol{y}_{j}^{\left(2\right)}$.
  6. 6.
    Each point $\boldsymbol{y}_{j}$ centres a trio and should be included in two other trios. Set the final weight for $\boldsymbol{y}_{j}$ to be half its weight in the trio it centres added to half the sum of its weights from the two other trios it is included in.
This approach approximates the curve geometry by fitting a circle to each set of three points; approximating the curve geometry is necessary to have an order of convergence greater than two; this approach is third-order. Note that this method also naturally handles multiple curves by only using nearest neighbours. If $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\tilde{N}_{\partial}}$ is dense enough on each constituent curve of $\partial S_{\pm}$, each trio of neighbouring points will belong to the same closed curve, correctly approximating the geometry of $\partial S_{\pm}$.
4 Numerical Tests
4.1 Genus-Two Surface
Implicit surfaces and surfaces of genus greater than zero can be challenging to mesh, particularly at low resolutions. Existing meshfree methods often assume that the distribution from which points are selected is known; without a parametrization of the surface, this distribution may be unknown, and it can be difficult to generate a uniform distribution.
For the next set of tests, we examine the surface $S$ defined as the zero level set of
(34) $\varphi\left(x,y,z\right):=\frac{1}{4\left(\left(x-1\right)^{2}+y^{2}\right)}+\frac{1}{4\left(\left(x+1\right)^{2}+y^{2}\right)}+\frac{1}{10}x^{2}+\frac{1}{4}y^{2}+z^{2}-1.$
This is a genus-two surface for which we previously computed Laplace-Beltrami eigenvalues in [30].
4.1.1 Average Values
Recall that Method 1 computes ratios of integrals: $\int_{S}f/\int_{S}g$. This means that it is straightforward to compute average values by setting $g=1$. To start, we use approximately uniformly generated point clouds using version B of algorithm A.1 of [30] with 25 test points per point in $S_{\tilde{N}}$. From a highly refined computation, we find that the average value of $x^{2}$ on the surface given by the zero set of $\varphi$ from (34) is 2.45884. We compare Method 1, using (33) to solve the Laplace-Beltrami problem, to integrating on a triangle mesh generated with the ball-and-pivot method [1] as implemented by the MATLAB Lidar Toolbox’s [27] $\mathtt{pc2surfacemesh}$, using linear interpolation between vertices (as in P1 finite elements). Integrating over the triangulation using a higher-order interpolation method would require a high-order approximation of the surface geometry and interpolation of additional points inside or outside the triangle (see [22] for example, for constructing a more accurate interpolant on a triangulation using RBFs).
For the Hilbert space $\mathcal{H}$, we use a separable basis as described at the end of Example 2.9 with
(35) $d_{x,n_{x}}=\left(\exp\left(q\sqrt{\frac{2\pi}{T}}\right)+\exp\left(q\sqrt{\left|\omega_{x,n_{x}}\right|}\right)\right)^{2}.$
$d_{y,n_{y}}$ and $d_{z,n_{z}}$ are defined similarly, substituting $\omega_{y,n_{y}}$ and $\omega_{z,n_{z}}$, respectively, in place of $\omega_{x,n_{x}}$. We use $T=12$, $\Omega=\left[-5,5\right]^{3}$ and $\left(2\cdot 80+1\right)^{3}$ Fourier basis functions (treated separably to form the $\boldsymbol{\Phi}$ matrix so that this is computationally feasible) to generate Figure 1, which shows the convergence of the meshfree method compared against triangulation and simply taking the average value on the point cloud (Point Cloud Average), as in a Monte Carlo method. The reference value we use to estimate the relative error is computed using a triangulation with 4.8 million vertices. We test a couple of different values of $q$; our observation is that for domains roughly of this size, a choice of $q$ between 2 and 4 works reasonably well. Functions with more oscillations are often better approximated using a lower value of $q$ to reduce the suppression of higher frequencies. However, as demonstrated in Figure 1, an optimal selection of $q$ is not necessary to achieve rapid convergence, and multiple values of $q$ could work well. It should be noted that the points are not quite uniformly generated, which explains why the “Point Cloud Average” method fails to converge. The point cloud with $\tilde{N}=800$ is shown in Figure 2 (Left).
Figure 1: Convergence plot for estimates of the average value of $x^{2}$ on a genus-two surface. Two shape parameters are shown for the meshfree method ($q=7/3,10/3$).
There are a few observations to make regarding Figure 1. For the triangulation approach, the error is high before falling into a pattern of consistent $\mathcal{O}\left(h_{\text{max}}^{2}\right)=\mathcal{O}\left(\tilde{N}^{-1}\right)$ convergence; this is due to meshing errors at low point resolutions. The meshfree methods reach a relative error on the order of $10^{-5}$ using only around two or three thousand points. This is roughly the error floor using the $\boldsymbol{\Phi}$ matrix, so other strategies (namely, complete orthogonal decomposition or singular value decomposition) would have to be employed to solve the optimization problem with higher accuracy. For a fast and reasonably accurate approach using a sparse point cloud, however, the separable basis method excels. It would take a mesh with nearly 100000 vertices to reach the level of accuracy achieved by the meshfree method with under 3000 points.
Also of note is that the meshfree approach does not require the point cloud to be “well-spaced”. That is, the point cloud may be significantly denser in some regions of the surface or have completely random point spacing, and the meshfree approach will still produce a reasonable result. Monte Carlo methods will clearly fail when points are drawn from an unknown, non-uniform distribution, and meshing algorithms tend to produce errors when the point cloud spacing is extremely uneven. We use randomly generated point clouds using algorithm A.1 of [30] (1 test point per point) for the tests in Table 1 and use $q=3$ for the meshfree method, with the other parameters left unchanged from Figure 1. Unlike in the previous test, the point cloud is not processed at all here after points are placed on the surface; the points are not as evenly spaced as in the previous test. This is shown in Figure 2 (Centre). Meshing these point clouds using MATLAB’s ball-and-pivot routine produces many missing elements, so we do not include it in this test.
Figure 2: Three point clouds on a genus-two surface with 800 points. Left: Regularly-spaced. Centre: Random. Right: Irregularly-spaced.
$\tilde{N}$ Relative Error (Random) Relative Error (Irregular)
400 4.5328E-02 5.0402E-01
800 5.6759E-03 2.0952E-01
1200 2.4342E-03 4.3325E-02
1600 6.1477E-04 3.5045E-02
2000 7.0428E-04 9.5554E-03
2400 2.3534E-05 2.0272E-03
Table 1: Left: Relative error for computing the average value of $x^{2}$ on a genus-two surface and using randomly generated points (Figure 2, Centre) and an uneven point distribution (Figure 2, Right). We use $q=3$ for both tests.
The error for the meshfree method is higher than that of using a more uniform point distribution as in Figure 1, but convergence still appears to be super-algebraic.
Finally, we produce highly non-uniform point clouds by oversampling points with $x$ values of small magnitude. Specifically, we place half of the points so that $x\in\left(-1,1\right)$, then sample the remaining points randomly over the whole surface. An example of such a point cloud is shown in Figure 2 (Right). Taking the average value of a function on the point cloud is clearly not an option for such a point distribution, and meshing such a point cloud will produce a large number of errors. The same parameters as the random point cloud test are used for Table 1 (Right), which shows the relative error for various values of $\tilde{N}$ on the irregular point cloud. The meshfree method improves consistently with each refinement. The errors are consistently higher here; this is not surprising, since $h_{\text{max}}$ is significantly larger with such an uneven point distribution.
4.1.2 Voronoi Domain Decomposition and Surface Area
Using the procedure of Subsection 3.5, we can divide the genus-two surface into a collection of subdomains, each diffeomorphic to a simply connected subset of the plane. An example of such a decomposition is shown in Figure 3 (Left), where both the points used for the Voronoi cells and the surface point cloud are generated using the same procedure as the first test in Subsection 4.1.1.
We compare the accuracy of Method 2 against triangulation in Table 2, where we compute the surface area of the genus-two surface. 100 surface patches are used for the meshfree approach, and the Hilbert space given by (35) is used, with $d_{n}$ replacing $d_{x,n_{x}}$ and $\left\|\boldsymbol{\omega}_{n}\right\|_{2}$ replacing $\left|\omega_{x,n_{x}}\right|$. The parameters $q=5,T=5$ are chosen, where each surface patch is scaled to fit in the cube $\left[-1/4,1/4\right]^{3}$, and then the computations for each patch are performed with a Hilbert space of functions on $\left[-1/2,1/2\right]^{3}$ so that the box side length of 1 is the same for each patch. We use $\left(2\cdot 11+1\right)^{3}$ Fourier basis functions in the box, and the resulting optimization problem ((33) leading to (16) when the basis is truncated). The constrained problem (16) is solved using a complete orthogonal decomposition (as opposed to solving $\boldsymbol{\Phi\beta}=\boldsymbol{f}$) for all remaining numerical tests.
Meshfree Triangulation
$\tilde{N}$ Estimate Rel. Diff. Estimate Rel. Diff.
4000 46.5490998641338 1.4987E-03 46.5787702744442 8.6147E-04
8000 46.6132362564690 1.2294E-04 46.5985576828876 4.3702E-04
16000 46.6184531258073 1.1038E-05 46.6088095697633 2.1712E-04
32000 46.6189600220989 1.6443E-07 46.6137557765005 1.1102E-04
64000 46.6189676876957 N/A 46.6163684686587 5.4973E-05
4800030 46.6189312401320 N/A
Table 2: Surface area estimates using the meshfree method and triangulation. Digits that match all subsequent estimates are coloured green, and the relative difference between the estimate and each method’s final estimate ($\tilde{N}=64000$ for the meshfree method, $\tilde{N}=4800030$ for the triangulation) is given as “Rel. Diff.”.
The rapid convergence of the meshfree method is apparent in this test; we find the first 6 or 7 digits of the surface area with $\tilde{N}=64000$ while only finding the first 4 with the triangulation. Also notable is that the meshfree estimate with $16000$ points appears to be significantly more accurate than the triangulation estimate with $64000$ vertices. The 4.8 million vertex triangulation is also likely less accurate than the meshfree estimates with 32000 or 64000 points.
Figure 3: Left: Voronoi cell subdomains on a genus-two surface. Right: Planar subdomain boundary for the genus-two surface.
4.1.3 Planar Boundary Surface Area Test
We repeat the test from 4.1.2 but with two subdomains created using a planar boundary as described in 3.5.2; the boundary curves are shown in Figure 3 (Right), discretized with 2000 points. The $z=0$ plane is used to separate the subdomains. The Hilbert space given by (35) is used again, with $d_{n}$ replacing $d_{x,n_{x}}$, $\left\|\boldsymbol{\omega}_{n}\right\|_{2}$ replacing $\left|\omega_{x,n_{x}}\right|$, and $q=5,T=10$, in a box with dimensions $10\times 6\times 2$ with $\left(2\cdot 13+1\right)^{3}$ Fourier modes. For this test, we use (32) to solve the Laplace-Beltrami problem. The surface area is estimated again in Table 3, with the relative error computed using the $\tilde{N}=64000$ result from Table 2 as an approximate “exact” value. We see that we are able to obtain an accurate estimate with far fewer points than the previous Voronoi test, though the Voronoi test used fewer points per subdomain, which is the main factor in computation time. This method achieves a more accurate estimate using $2560$ points than a triangulation does with 64000, and the 2560 point result is competitive with the 4.8 million vertex triangulation. We also estimate the convergence rate in Table 3 and see that the method appears to converge at a high-order rate as expected.
$\tilde{N}$ Estimate Relative Error Convergence ($\mathcal{O}\left(h_{\text{max}}^{p}\right)$)
320 41.9984565553679 9.9112E-02 N/A
640 45.6178349766375 2.1475E-02 4.413
1280 46.5745212157448 9.5340E-04 8.987
2560 46.6180662106845 1.9337E-05 11.247
5120 46.6189398270088 5.9763E-07 10.032
Table 3: Surface area estimates using the meshfree method with a planar boundary between subdomains. Convergence rate is relative to $h_{\text{max}}$, which scales as $\mathcal{O}\left(\tilde{N}^{-\frac{1}{2}}\right)$.
4.2 High-Order Approximation of Singular Integrals
In certain applications, particularly in the numerical approximation of PDEs, the function to be integrated has a singularity. Crucially, the location of the singularity is typically known. As it turns out, meshfree methods are well suited to approximating these integrals with a slight modification of the methods described in Subsection 2.1.
Assume that we want to integrate a function $f:S\to\mathbb{R}$ that has a singularity at some point $\boldsymbol{x}_{0}\in S$. As before, we could attempt to find a function $u$ such that $\Delta_{S}u=f$ to convert $\int_{S}f$ to an integral on $\partial S$. The issue we immediately encounter is that $u$ may also have a singularity at $\boldsymbol{x}_{0}$, which would mean that $u$ would not be in any of the Hilbert spaces of functions commonly used for meshfree methods, which typically consist of functions that are at least continuous. Furthermore, we still need $\Delta_{S}v$ to be bounded on $S_{\tilde{N}}$ for any $v$ in the Hilbert space in order to ensure boundedness of the operator $\mathcal{L}$ from Subsection 2.1. This means that the functions in $\mathcal{H}$ need to be at least $C^{2}$ at every point in $S_{\tilde{N}}$. That is, we need to search for functions that are possibly singular at $\boldsymbol{x}_{0}$, but are $C^{2}$ everywhere else; such functions cannot exist in the Hilbert spaces we have used so far.
Therefore, we need to alter our framework slightly. Rather than a Hilbert space of functions, we consider a Hilbert space $\mathcal{H}$ equipped with a (not necessarily injective) linear evaluation operator $\mathcal{E}$ and consider a problem analogous to (5):
(36) $\text{minimize, over }a\in\mathcal{H}:$ $\left\|a\right\|_{\mathcal{H}},$
$\text{subject to}:$ $\left(\mathcal{F}\left(\mathcal{E}a\right)\right)\left(\boldsymbol{x}_{j}\right)=f\left(\boldsymbol{x}_{j}\right)\text{ for each $j$$\in\left\{1,2,\ldots,\tilde{N}\right\},$}$
$\left(\mathcal{G}\left(\mathcal{E}a\right)\right)\left(\boldsymbol{y}_{j}\right)=g\left(\boldsymbol{y}_{j}\right)\text{ for each $j$$\in\left\{1,2,\ldots,\tilde{N}_{\partial}\right\}.$}$
$\mathcal{E}$ in this case must map $a\in\mathcal{H}$ to a function that is $C^{p}$, where $\mathcal{F},\mathcal{G}$ have a maximum order of $p$, except possibly at a finite number of points in $S$ (but not in $S_{\tilde{N}}$ or $\partial S_{\tilde{N}_{\partial}}$). With a suitable choice of $\mathcal{H}$ and $\mathcal{E}$ it is then possible to ensure that the constraint set for (36) is non-empty and closed. If a solution in $a\in\mathcal{H}$ exists such that $\mathcal{F}\left(\mathcal{E}a\right)=f$ on $S$ and $\mathcal{G}\left(\mathcal{E}a\right)=g$ on $\partial S$, convergence is guaranteed due to Proposition 2.1. We illustrate this approach with a couple of examples.
4.2.1 Singular Integrand in 2D
Integrals of the form $\int_{S}f\left(\boldsymbol{y}\right)K\left(\boldsymbol{x},\boldsymbol{y}\right)\,\text{d}\boldsymbol{y}$, where $K$ has a singularity when $\boldsymbol{x}=\boldsymbol{y}$, are common when solving PDEs by boundary integral approaches and appear in various physical applications, particularly in electromagnetism. Let $S$ be the unit disk and then consider the example:
$\frac{1}{2\pi}\int_{S}\ln\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}\,\text{d}\boldsymbol{x}=\begin{cases}\frac{1}{4}-\frac{\left\|\boldsymbol{x}_{0}\right\|_{2}^{2}}{4},&\left\|\boldsymbol{x}_{0}\right\|_{2}\leq 1\\ -\frac{1}{2}\ln\left\|\boldsymbol{x}_{0}\right\|_{2},&\left\|\boldsymbol{x}_{0}\right\|_{2}>1\end{cases}.$
Physically, this corresponds to the electric potential inside and outside an infinitely long cylindrical wire with constant charge density, up to a constant; the result as a function of $\boldsymbol{x}_{0}$ solves $\Delta u=1$ in $S$ and 0 outside $S$. To integrate this, we search for a function $u$ such that $\Delta u=-\frac{1}{2\pi}\ln\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}.$
One possible function $u$ is $\frac{1}{8\pi}\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{2}\ln\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}+\frac{1}{8\pi}\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{2}$; we could simply use this to convert the integral into a boundary integral, but of course the goal is to be able to integrate functions when we do not have an expression for $u$. For example, if $S$ were instead a curved surface and $\Delta u$ were swapped for $\Delta_{S}u$, we would not have such an expression for a possible $u$ available in general.
Returning to the optimization problem from (36), for this problem, we will need to solve
(37) $\text{minimize, over }a\in\mathcal{H}:$ $\left\|a\right\|_{\mathcal{H}},$
$\text{subject to}:$ $\left(\Delta\left(\mathcal{E}a\right)\right)\left(\boldsymbol{x}_{j}\right)=-\frac{1}{2\pi}\ln\left\|\boldsymbol{x}_{j}-\boldsymbol{x}_{0}\right\|_{2}$
for each $j\in\left\{1,2,\ldots,\tilde{N}\right\}$,
When $\boldsymbol{x}_{0}\in S$, $\frac{1}{2\pi}\ln\left\|\boldsymbol{x}_{j}-\boldsymbol{x}_{0}\right\|_{2}$ is singular at $\boldsymbol{x}_{0}$, and therefore requires our operator $\mathcal{E}$ to take elements of $\mathcal{H}$ to a function with a singularity at $\boldsymbol{x}_{0}$. To be more concrete, let $\mathcal{H}=\ell^{2}\times\ell^{2}$, then, where $\Omega$ is again a box that contains $S$, let $\mathcal{E}:\mathcal{H}\to C\left(\Omega\right)\cap C^{2}\left(\Omega\setminus\left\{\boldsymbol{x}_{0}\right\}\right)$ be such that
$\mathcal{E}a=u+sv,$
where $u,v\in C^{q}\left(\Omega\right)$ vary with $a$ and $s\in C^{2}\left(\Omega\setminus\left\{\boldsymbol{x}_{0}\right\}\right)$ is fixed and $\Delta s$ has terms that capture the expected singularity. For example, $\Delta\left(u+sv\right)=-\frac{1}{2\pi}\ln\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}$ is solved by $u\left(\boldsymbol{x}\right)=\frac{1}{4\pi}\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{2},v\left(\boldsymbol{x}\right)=-\frac{1}{16\pi},s\left(\boldsymbol{x}\right)=\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{2}\ln\left(\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{2}\right)$. Assuming we did not know this exact solution, but at least knew $\Delta\left(u+sv\right)$ needed a logarithm singularity with the scaling and any non-singular terms undetermined, we could use $s\left(\boldsymbol{x}\right)=\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{2}\ln\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{2}$ and
(38) $\left(\mathcal{E}\left(a,b\right)\right)\left(x\right):=\left(\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}a_{n}e^{i\boldsymbol{\omega}_{n}\cdot\boldsymbol{x}}\right)+s\left(\boldsymbol{x}\right)\left(\sum_{n=1}^{\infty}\tilde{d}_{n}^{-\frac{1}{2}}b_{n}e^{i\tilde{\boldsymbol{\omega}}_{n}\cdot\boldsymbol{x}}\right),$
where $\left\{\boldsymbol{\omega}_{n}\right\}$ and $\left\{\tilde{\boldsymbol{\omega}}_{n}\right\}$ correspond to Fourier frequencies on the box $\Omega$. Note that $d_{n}$ and $\tilde{d}_{n}$ control the smoothness of $u$ and $v$, respectively.
It is important to note that $\Delta s,\partial_{x}s,$ and $\partial_{y}s$ should not be zero on $\Omega\setminus\left\{\boldsymbol{x}_{0}\right\}$ with a delta distribution singularity at $\boldsymbol{x}$; this would allow arbitrary multiples of $s,xs$ or $ys$ to be added to our solution and change the resulting integral (due to the delta distribution).
For our chosen $s$, a logarithm term appears in $\Delta s$, which will allow $\ln\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}$ to be integrated. After computing $u+sv$ so that $\Delta\left(u+sv\right)=-\frac{1}{2\pi}\ln\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}$ on $\tilde{N}$ scattered points, the boundary integral $\oint_{\partial S}\nabla\left(u+sv\right)\cdot\hat{\boldsymbol{n}}_{\partial S}$ is computed on 1000 evenly spaced points on the unit circle with the standard trapezoid rule (which converges super-algebraically for $C^{\infty}$ functions for the unit circle). We use $d_{n}=\left(\exp\left(q\sqrt{2\pi/T}\right)+\exp\left(q\sqrt{\left\|\boldsymbol{\omega}_{n}\right\|_{2}}\right)\right)^{2},q=4,T=10$, $\Omega=\left[-2,2\right]^{2}$, and $\left(2\cdot 30+1\right)^{2}$ Fourier basis functions. Interior points are generated using algorithm A.2 of [30] with 20 test points per point and no additional weighting of points near the boundary. The relative error for computing $\frac{1}{2\pi}\int_{S}\ln\left\|\boldsymbol{x}\right\|_{2}\,\text{d}\boldsymbol{x}=\frac{1}{4}$ is shown in Figure 4. High-order convergence is observed, where we note that the point spacing varies as $h_{\text{max}}\propto\tilde{N}^{-\frac{1}{2}}$. We note again that a unit circle is only used as a test problem to compare to a known solution; the method’s main use is for irregularly-shaped domains and surfaces.
Figure 4: Relative error for computing $\frac{1}{2\pi}\int_{S}\ln\left\|\boldsymbol{x}\right\|_{2}\,\text{d}\boldsymbol{x}$ on $\tilde{N}$ scattered points on the unit disk.
4.2.2 Singular Integrand on a Surface
We now consider the surface described by $S:\left\{\left(x,y,z\right):\varphi\left(x,y,z\right)=0,z\geq 0\right\}$ where $\varphi$ is the level set for a paraboloid: $\varphi\left(x,y,z\right):=z+x^{2}+y^{2}-1.$ We compute
(39) $\frac{1}{4\pi}\int_{S}\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{-1}\,\text{d}\boldsymbol{x},$
where $\boldsymbol{x}_{0}=\left(0,0,1\right)\in S$. This integral can be interpreted as the electric potential at $\boldsymbol{x}_{0}$ produced by a constant surface charge on $S$. We integrate by attempting to find a function of the form $u+sv$ such that $\Delta_{S}\left(u+sv\right)=-\frac{1}{4\pi}\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{-1}$, where $\Delta_{S}$ is now the Laplace-Beltrami operator, and then converting the surface integral to a line integral on the unit circle in the $xy$-plane. The line integral is again evaluated using the trapezoid rule and 1000 evenly spaced points. The Laplace-Beltrami problem is discretized using a method similar to the one used in the previous test (see Eq. (37)), but with $\Delta_{S}$ computed using (31) replacing $\Delta$. We seek a solution of the form $u+sv$ where
$s\left(\boldsymbol{x}\right)=\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2},$
and $u,v\in C^{\infty}\left(\Omega\right)$ are unknown. This choice of $s$ gives $\Delta s=2\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{-1}$; however, $\Delta s$ is only one term in $\Delta_{S}s$. The necessary functions $u$ and $v$ to solve $\Delta_{S}\left(u+sv\right)=-\frac{1}{4\pi}\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{-1}$ will vary depending on the local geometry of the surface and will be computed numerically. The idea is therefore to use a known Poisson solution in 3D space to handle the desired singular behaviour in the surface Poisson problem.
The Hilbert space and $\mathcal{E}$ used are identical to Subsection 4.2.1, but with $\Omega=\left[-2,2\right]^{3}$ (instead of $\left[-2,2\right]^{2}$) and $\left(2\cdot 11+1\right)^{3}$ Fourier basis functions. Surface points are generated using version B of algorithm A.1 of [30] with 25 test points per point; notably, we do not place extra points near the singularity. This is compared to the naive approach, where a solution $u\in C^{\infty}\left(\Omega\right)$ is sought, in Table 4 (effectively searching for solutions of the form $\mathcal{E}\left(a,0\right)$ only, where $\mathcal{E}$ is similar to (38), but for 3 dimensions). We compare to a value computed to machine precision from a 1D integral resulting from a parametrization, and see that the augmented method converges to the first 9 digits by $\tilde{N}=2560$, while the naive approach fails to converge beyond a digit or two. The Poisson solution using the augmented approach for $\tilde{N}=2560$ is shown in Figure 5.
Augmented Naive
$\tilde{N}$ Estimate Relative Error Estimate Relative Error
80 -0.633973868 1.3666E-04 -0.699667536 1.0347E-01
160 -0.634021133 6.2115E-05 -0.662463834 4.4796E-02
320 -0.634028690 5.0196E-05 -0.637664974 5.6847E-03
640 -0.634060749 3.6451E-07 -0.491952726 2.2412E-01
1280 -0.634060516 1.6639E-09 -0.349696715 4.4848E-01
2560 -0.634060518 1.9579E-11 -0.554560171 1.2538E-01
Table 4: Estimate of (39) using solutions augmented with a singular term (Augmented) and a naive approach (Naive).
Figure 5: A computed solution to $\Delta_{S}u=-\frac{1}{4\pi}\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{-1}$ on the paraboloid using the augmented method with $\tilde{N}=2560$ points ($\boldsymbol{x}_{0}=\left(0,0,1\right)$).
5 Conclusions
We presented, analyzed, and tested two high-order, meshfree techniques for surface integration. The first method (Subsection 3.1) uses a result relating solvability and boundedness (Proposition 2.1) to determine which scaling of a function makes a Poisson problem solvable, then deduces the ratio of the integral of two functions on a domain of interest. This approach is useful for finding the average values of functions on surfaces, as shown in the tests of 4.1.1. We also demonstrated numerically that reasonable estimates of the average value of a function can be obtained using our approach even on irregular, non-uniform point distributions. The second method (Subsection 3.2) uses the divergence theorem to reduce surface integrals to line integrals and takes advantage of the ability of symmetric meshfree methods to find one of many possible solutions to PDEs; this avoids imposing boundary conditions on boundaries that are merely piecewise-smooth. We demonstrated method 2’s accuracy in 4.1.2 and 4.1.3, where we estimated the surface area of a genus-two surface considerably more accurately than a standard triangulation approach. Both methods also preserve the super-algebraic approximation properties of meshfree interpolation methods and do not require prior knowledge of the probability distribution from which a point cloud is sampled; this differs from Monte Carlo methods which are $\mathcal{O}\left(\tilde{N}^{-\frac{1}{2}}\right)$ and typically assume a certain sampling distribution.
In Subsection 4.2, we generalized the optimized Hermite-Birkhoff problem solved by symmetric meshfree methods to handle singularities while maintaining high-order convergence. We then used the approach for solving a Poisson problem with a singular source term to evaluate singular integrals in 2D and on a surface, achieving rapid convergence in both cases.
Additional concepts we introduced included separable basis functions for Fourier series meshfree methods (see the end of Example 2.9), which allow for the rapid and systematic evaluation of matrices that appear when using a symmetric matrix approach to solving the optimization problem; solving $\boldsymbol{\Phi\beta}=\boldsymbol{f}$ directly is known as RBF-direct in the literature when radial basis functions are used. Constructing the correct $\boldsymbol{\Phi}$ matrices when using Hermite RBFs can be difficult, particularly for surface PDEs, since it involves computing multiple derivatives of the original RBF. Even for a simple Laplace-Beltrami problem, all fourth-order derivatives of the RBF must be computed to form the correct $\boldsymbol{\Phi}$, which can be cumbersome. Fourier series, particularly separable ones, are straightforward to work with by comparison and allow for greater flexibility in balancing conditioning and convergence, since the weights $d_{n}>0$ can be chosen freely. We also introduced simple, automated approaches for implicit surface domain decomposition for meshfree methods (Subsection 3.5).
In future work, we seek to use the introduced integration techniques for solving PDEs numerically (in weak form or via boundary integral methods). This will involve developing a procedure to “stabilize” the integration weights; high-order integration schemes, including the ones presented here, can produce negative integration weights. These jeopardize the stability of time-stepping methods. We have observed that points with negative weights can be repeatedly removed to eventually produce a positive set of weights while maintaining accuracy, but this warrants further testing. Adaptive point selection techniques are another area of future work, since the meshfree nature of these approaches allows complete freedom in point placement without the need to repeatedly re-mesh.